OpenStudy (anonymous):

Given: 3x + y = 1. Solve for y. y = -3x + 1 y = -3x - 1 y = 3x - 1

OpenStudy (istim):

Rearrange for y, basically.

OpenStudy (anonymous):

y = -3x - 1

OpenStudy (istim):

Or, let Eric give the answer. Corey will totally learn this way.

OpenStudy (anonymous):

step 1. add -3x to both sides.. and you'll get: \[\LARGE \underbrace{3x-3x}_{\text{This is 0 }}+y=1-3x\] so... \[\LARGE \cancel{3x}-\cancel{3x}+y=1-3x\] so what do you have?

OpenStudy (anonymous):

@Eric_Playsinacoolway that's not correct.

OpenStudy (anonymous):

sorry! didn't follow the formula @Kreshnik

OpenStudy (istim):

No, Corey.

OpenStudy (anonymous):

so then it is the first one

OpenStudy (anonymous):

You tell us.

OpenStudy (istim):

Kreshnik has it perfectly laid out, actually.

OpenStudy (anonymous):

@CoreyContini -3+1=1-3 you can chenge places... and \[\LARGE y=1-3x\] what can you get?? :)

OpenStudy (anonymous):

@IsTim thank you ^^

OpenStudy (istim):

Yup.